# Towards grid-aware multi-period flexibility aggregation – A constrained zonotope approach

### v1

@@ Abstract
Aggregation schemes provide a means to reduce the computational complexity of power system operation by reducing the number of devices that are considered individually. This can be achieved with tools of computational geometry, where the feasible set is projected onto the decision variables of the point of interconnection. Set projection is computationally expensive, especially in the context of multi-period power system operation. Hence this calls for efficiency improvements via structure exploitation of certain set representations, such as constrained zonotopes. This paper proposes these benefits for efficient flexibility aggregation. We evaluate the performance of the proposed method on a 15-bus distribution grid with time-dependent elements for up to 96 timesteps. The results suggest that the presented method significantly improves computation times.
@@ I Introduction
To address these challenges, hierarchical optimization has been widely adopted in the literature. In the first stage of the approach, flexibility aggregation, the Feasible Operation Region (FOR) is computed. The FOR describes the values of the shared variables, usually active and reactive power, at the interconnection point that can be achieved without compromising the feasibility of the optimization problem from the DSO’s perspective. A survey of different methods can be found in [1]. The TSO can then use this information to determine the optimal utilization of DER flexibility without requiring direct data from the distributed grid assets.
@@
The standard FOR estimation methods rely on sampling the region and computing the convex hull of these points. The samples can be randomly obtained [2] or calculated in a more structured manner, utilizing knowledge of the grid and constraints to solve optimization problems [3, 4, 5]. The methods above consider assumptions that can be quite limiting. They assume a unique interconnection point between DSO and TSO, the voltage at this point to be constant, and no time-dependent elements, e.g., storage systems.
@@
There is limited research on the flexibility aggregation in distribution grids that considers time-dependent elements. A survey of various methods for aggregating batteries can be found in [6]. They offer computational tractability but neglect grid constraints or rely on data-driven formulations, which reduces their accuracy for grid-aware applications. Multi-period cost curve formulations [7] capture temporal coupling but limits the FOR to only portraying active power.
@@
Recently, set projection has been used for the computation of the FOR [8, 9]. The method has been extended to FORs, which include time-dependent elements [10], but it requires solving multiple optimization problems. For affine grid models, such as LinDistFlow [11], the projection can be obtained in one step via tools of computational geometry. The downside of this technique is the computational burden of the set projection operation, which limits its scalability [12].
@@
One line of research addresses this computational burden of polytope projection by applying zonotope approximations [13, 14]. However, such approximations are not always accurate, particularly when the underlying set is non-symmetric.
@@
The present paper introduces a novel approach to compute time-coupled FORs by extending the non-iterative projection-based formulation [15] to multi-period settings. To overcome computational bottlenecks that normally hinder the scalability of set projections, we rely on Constrained Zonotopes ($\mathcal{CZ}$[16]. CZs enable non-symmetric set representation with efficient set projection properties. This results in significant reduction of computation times without compromising representation accuracy. Furthermore, most of the remaining computation can be carried out offline, enabling the practical implementation for the computation of FORs for a 15-bus distribution grid that include time-dependent elements such as storage units for up to 96 timesteps.
@@
The remainder of this paper is structured as follows: Section II elaborates the problem statement and recaps the flexibility aggregation in the ADP approach. In Section III we present how $\mathcal{CZ}\mathrm{s}$ can be utilized for efficient flexibility aggregation over time, while Section IV highlights the potential of this method for dynamic problem settings. Section V compares and discusses numerical results for the presented methods. The paper concludes with Section VI.
@@ II Problem Statement
We define the nodal active and reactive power $p_{m}(k)$ and $q_{m}(k)$ at each timestep $k$ in the considered interval $\mathcal{I}={1,\dots,N}$ as
$p_{m}(k)=\sum_{l\in\mathcal{N_{d}}}p_{m,l}(k),\qquad q_{m}=\sum_{l\in\mathcal{N}}q_{m,l}(k),$ (1)
and the net powers at each node
@@
We limit the reactive power support from each renewable generator by constraining the maximum apparent power $\bar{s}_{m}(k)$ and the power factor limit $\alpha$ at each timestep with the affine constraints [17]
$p_{m}^{g}(k)\leq\bar{s}_{m}(k)\cos(\alpha),\quad-p_{m}(k)\leq\alpha q_{m}(k)\leq p_{m}(k).$ (4)
Next, we formulate the active/reactive power flow $p_{m,l}(k)$ and $q_{m,l}(k)$ on each line $(m,l)\in\mathcal{B}_{d}$ and the squared voltage $\nu_{m}$ at each bus $m\in\mathcal{N}_{d}$ according to the DistFlow formulation for radial grids [18]
@@
where, $r_{m,l}$ and $x_{m,l}$ are, respectively, the line reactance and capacitance and $\ell_{m,l}(k)$ is the squared line current.
To obtain a polyhedral constraint set, we use the first-order Taylor series as linear approximation for the square current $\ell_{m,l}(k)$ around the operational point $t_{m,l}^{0}$ [19],
$\ell_{m,l}(k)\approx\ell_{m,l}(k)^{0}+J_{m,l}|_{\begin{subarray}{c}t_{m,l}^{0}(k)\end{subarray}}^{\intercal}\delta_{m,l}(k)$ (6)
@@
$J_{m,l}|_{\begin{subarray}{c}t_{m,l}^{0}\end{subarray}}$ $=\begin{bmatrix}\frac{2p_{m,l}(k)}{\nu_{l}^{0}}\\ \frac{2q_{m,l}(k)}{\nu_{l}^{0}}\\ -\frac{(p_{m,l}^{0})^{2}+q_{m,l}^{0})^{2}}{(\nu_{l}^{0})^{2}}\end{bmatrix}$. (8)
Additionally, we limit the squared voltage and the squared current
@@
Moreover, let $\mathcal{S}\subseteq\mathcal{N}$ denote the set of buses equipped with storage units. Each storage unit has a limited energy capacity $e_{m}^{s}(k)$, bounded by:
@@
with the maximum charging and discharging rate of
@@
In this work we neglect the reactive power support of storage systems.
A simple discrete-time integrator dynamic model determines the State of Charge (SOC) of the storage unit. Assuming a fixed timestep size $\Delta k$ and neglecting losses, the SOC dynamic can be described as
$e_{m}(k+1)=e_{m}(k)-\Delta k\cdot p_{m}^{s}(k),\quad m\in\mathcal{S}.$ (12)
The initial condition of the battery is given by
$e_{m}(1)=e_{m}^{0},\quad m\in\mathcal{S}.$ (13)
We define the resulting constraint set as
@@
Figure 1: Tree-structured optimization of power system operation adapted from [15].
The resulting constraint set is a convex and bounded polyhedron, i.e. a polytope. The set projection of $\mathcal{X}\subseteq\mathcal{Y}\times\mathcal{Z}$ onto $\mathcal{Z}$ is defined as in [20]
$\operatorname{proj}_{\mathcal{Z}}(\mathcal{X})\doteq\{z\in\mathcal{Z}\ |\ \exists y\in\mathcal{Y}\text{ with }(z,y)\in\mathcal{X}\}\subseteq\mathcal{Z},$ (14)
which gives us the possible values for the coupling variables that can be achieved without violating any of the constraints of the original set, i.e., we can compute the FOR $\mathcal{F}_{i}$ as
@@ III-A Approximation of Feasible Operating Region
with a constraint matrix $A\in\mathbb{R}^{m\times n}$ and a constraint vector $b\in\mathbb{R}^{m}$. This formulation enables the projection of the feasible set onto coupling variables using methods from computational geometry, such as Fourier–Motzkin elimination or vertex enumeration [21, pp. 2-3]. However, for multi-stage flexibility aggregation with storage units, the increased problem dimension renders these algorithms computationally intractable.
@@ III-B Reformulation of the Feasible Set
with a center vector $c\in\mathbb{R}^{n}$, a generator matrix $G\in\mathbb{R}^{n\times n_{g}}$ and the scalar internal factors $\alpha_{i}$. Zonotope projection is performed via a simple sparse linear map and therefore can be computed very efficiently. In the following, the shorthand notation $\mathcal{Z}=\langle c,G\rangle_{\mathcal{Z}}$ is used for zonotopes. However, zonotopes are limited to model centrally symmetric convex polytopes. To address this limitation Constrained Zonotopes ($\mathcal{CZ}\mathrm{s}$) extend zonotopes by adding linear constraints to the internal factors $\alpha_{i}$ [16]. This results in the definition
@@
with center vector $c\in\mathbb{R}^{n}$, generator matrix $G\in\mathbb{R}^{n\times n_{g}}$, constraint matrix $A\in\mathbb{R}^{m\times n_{g}}$, and constraint vector $b\in\mathbb{R}^{m}$. In the following, the shorthand notation $\mathcal{CZ}=\langle c,G,A,b\rangle_{\mathcal{CZ}}$ is used for constrained zonotopes. Without limitations on the number of generators $n_{g}$ and the number of constraints $m$, $\mathcal{CZ}\mathrm{s}$ can express arbitrary convex polytopes and offer many of the computational advantages of zonotopes, like projection via matrix multiplication [16].
@@
Given a feasible set as convex and bounded polytope one can determine upper and lower bounds for each variable by solving two Linear Programms (LPs) for each variable $x_{i,k}$. This allows to determine the vectors of upper and lower variable bounds $ub_{x}$ and $lb_{x}$, which define the bounding zonotope $Z_{b}$ with center $c$ and generator matrix $G$ that over-approximates the given polytope $Z_{b}=\langle c,G\rangle_{Z}\subseteq\mathcal{P}$ as follows:
@@
However, as shown in Subsection V-B selecting the variable bounds sufficiently large is adequate without loss of accuracy for any $\tilde{lb}$ and $\tilde{ub}$ as long as
@@
Given the feasible set $\mathcal{X}_{i}$ of the multi-stage OPF as polytope $\mathcal{P}$ in H-rep, it is defined as an intersection of a finite number of halfspaces $H_{i}=\{x\in\mathbb{R}^{n}\mid h^{\top}x\leq\zeta\}$. To reconstruct the polytope as $\mathcal{CZ}$, the bounding zonotope is successively intersected with each halfspace $H_{i}$. Based on the method in [22], each intersection updates the $\mathcal{CZ}$ as follows:
@@
This step assumes that the provided polytope is irredundant, so that no feasibility checks are required before intersection. Otherwise, it is beneficial to keep the representation compact by solving an LP to check whether a halfspace intersects the current $\mathcal{CZ}$ and therefore has to be added to the representation [22]. The proposed approach is summarized in Algorithm 1.
@@
1:for all $H_{i}$ modeling static constraints (5)-(2), (4), (9) do
@@
1:for all $H_{i}$ modeling dynamic constraints like (3) do
@@
Return $\mathcal{CZ}$
@@
Finally, once the $\mathcal{CZ}$ is constructed, projection onto the desired subspace is performed via the sparse linear map $M_{p}$, yielding:
@@
This projection operation utilizes the computational benefits of the $\mathcal{CZ}$ representation, enabling fast and scalable dimensionality reduction via linear map.
@@ IV Intersection with Additional Linear Constraints
For more dynamic use cases, it is essential to limit the computation cost required to consider changes in the feasible set $\mathcal{X}_{i}$ due to control inputs. Given the inherent separation of the set reformulation from the projection calculation in the presented method, most of the computation cost can be performed as an offline computation. Further, we can utilize that $\mathcal{CZ}\mathrm{s}$ allow to calculate halfspace and hyperplane intersections. This allows to add additional constraints to the $\mathcal{CZ}$ after initial setup, which is especially beneficial for increasing problem size, as most of the problem remains identical per time step. The static constraints (5)-(2), (4), (9) in $\mathcal{X}_{i}$ result in most of the offline computation time, which can be performed once. Then, the resulting $\mathcal{CZ}$ is modified as needed with reduced computation time, i.e. setpoint constraints can be adapted dynamically by adding the respective constraints via intersection each iteration. For instance, this enables the integration of revised forecast data in (3). The computational efficiency of these intersections is evaluated in Subsection V-B.

### v2

@@ Abstract
Aggregation schemes provide a means to reduce the computational complexity of power system operation by reducing the number of devices that are considered individually. This can be achieved with tools of computational geometry, where the feasible set is projected onto the decision variables of the point of interconnection. Set projection is computationally expensive, especially in the context of multi-period power system operation. This calls for efficiency improvements via structure exploitation of set representations. This paper proposes efficient flexibility aggregation via constrained zonotopes. We evaluate the performance of the proposed method on a 15-bus distribution grid with time-dependent elements for up to 96 timesteps. The results suggest that the presented method significantly improves computation times compared to classic polytope projection approaches.
[…]
TF and CR acknowledge financial support from TRR 391 Spatio-temporal Statistics for the Transition of Energy and Transport (520388526) by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation).
Submitted to the 24th Power Systems Computation Conference (PSCC 2026).
@@ I Introduction
To address these challenges, hierarchical optimization has been widely adopted in the literature. In the first stage of the approach, flexibility aggregation, the Feasible Operation Region (FOR) is computed. The FOR describes the values of the shared variables, usually active and reactive power, at the interconnection point that can be achieved without compromising the feasibility of the optimization problem from the DSO’s perspective. A survey of different methods can be found in [undef]. The TSO can then use this information to determine the optimal utilization of DER flexibility without requiring direct data from the distributed grid assets.
@@
The standard FOR estimation methods rely on sampling the region and computing the convex hull of these points. The samples can be randomly obtained [undefa] or calculated in a more structured manner, utilizing knowledge of the grid and constraints to solve optimization problems [undefb, undefc, undefd]. The methods above consider assumptions that can be quite limiting. They assume a unique interconnection point between DSO and TSO, the voltage at this point to be constant, and no time-dependent elements, e.g., storage systems. To mitigate some of these limitations, recent optimization-bases approaches compute inner-approximations of the aggregated flexibility in multi-period settings [undefe, undeff]. Similar approaches extend this approach to consider uncertainties [undefg] or solve optimization problems to parameterize a virtual battery model considering time-dependent elements and network constraints [undefh], but these inner approximation based methods remain inherently constrained by the geometry of the chosen surrogates.
@@
There is limited research on the flexibility aggregation in distribution grids that considers time-dependent elements. A survey of various methods for aggregating batteries can be found in [undefi, undefj]. They offer computational tractability but neglect grid constraints or rely on data-driven formulations, which reduces their accuracy for grid-aware applications. Multi-period cost curve formulations [undefk] capture temporal coupling but limits the FOR to only portraying active power.
@@
Recently, set projection has been used for the computation of the FOR [undefh, undefl]. The method has been extended to FORs, which include time-dependent elements [undefm], but it requires solving multiple optimization problems. For affine grid models, such as LinDistFlow [undefn], the projection can be obtained in one step via tools of computational geometry. The downside of this technique is the computational burden of the set projection operation, which limits its scalability [undefo].
@@
One line of research addresses this computational burden of polytope projection by applying zonotope approximations [undefp, undefq]. However, such approximations are not always accurate, particularly when the underlying set is non-symmetric.
@@
The present paper introduces a novel approach to compute time-coupled and voltage dependent active and reactive power FORs by extending the non-iterative projection-based formulation [undefr] to multi-period settings with time-dependent elements. To overcome computational bottlenecks that normally hinder the scalability of set projections, we rely on Constrained Zonotopes ($\mathcal{CZ}$[undefs]. $\mathcal{CZ}$s enable non-symmetric set representation with efficient set projection properties. This results in significant reduction of computation times without compromising representation accuracy. Furthermore, most of the remaining computation can be carried out offline, enabling the practical implementation for the computation of FORs for a 15-bus distribution grid that include time-dependent elements such as storage units for up to 96 timesteps.
@@
The remainder of this paper is structured as follows: Section II elaborates the problem statement and recaps the flexibility aggregation via polytope projection. In Section III we present how $\mathcal{CZ}\mathrm{s}$ can be utilized for efficient flexibility aggregation over time, while Section IV highlights the potential of this method for dynamic problem settings. Section V compares and discusses numerical results for the presented methods. The paper concludes with Section VI.
@@ II Problem Statement
We define the nodal active and reactive power $p_{m}(k)$ and $q_{m}(k)$ of node $m$ at each timestep $k$ in the considered interval $\mathcal{I}={1,\dots,N}$ as
$p_{m}(k)=\sum_{l\in\mathcal{N}_{d}}p_{m,l}(k),\qquad q_{m}=\sum_{l\in\mathcal{N}_{d}}q_{m,l}(k),$ (1)
Figure 1: Tree-structured optimization of power system operation adapted from [undefr].
with $p_{m,l}(k)$ and $q_{m,l}(k)$ representing the active and reactive power flow between nodes $m$ and $l$ at each timestep $k$. The net powers at each node are
@@
We limit the reactive power support from each renewable generator by constraining the maximum apparent power $\bar{s}_{m}(k)$ and the power factor limit $\cos(\phi)$ at each timestep with the affine constraints [undeft]
$p_{m}^{g}(k)$ $\leq\bar{s}_{m}(k)\cos(\phi),$ (4a)
$-p_{m}^{g}(k)\tan(\phi)$ $\leq q_{m}^{g}(k)\leq p_{m}^{g}(k)\tan(\phi).$ (4b)
Next, we formulate the active/reactive power flow $p_{m,l}(k)$ and $q_{m,l}(k)$ on each line $(m,l)\in\mathcal{B}_{d}$ and the squared voltage $\nu_{m}$ at each bus $m\in\mathcal{N}_{d}$ according to the DistFlow formulation for radial grids [undefu]
@@
where, $r_{m,l}$ and $x_{m,l}$ are, respectively, the line resistance and reactance and $\ell_{m,l}(k)$ is the squared line current.
To obtain a polyhedral constraint set, it is common to neglect line impedances. Then the squared branch current $\ell_{m,l}(k)$ does not need to be calculated. As shown in [undefr], this results in a relatively good approximation of the true flexibility region with a slight offset due to the neglected losses. Similar to [undefv, undefw, undefx] we use the first-order Taylor series as linear approximation for the square current $\ell_{m,l}(k)$ around the current operational point $t_{m,l}^{0}$ [undefw],
$\ell_{m,l}(k)\approx\ell_{m,l}(k)^{0}+J_{m,l}^{\top}|_{\begin{subarray}{c}t_{m,l}^{0}(k)\end{subarray}}\delta_{m,l}(k)$ (6)
@@
$J_{m,l}|_{\begin{subarray}{c}t_{m,l}^{0}\end{subarray}}$ $=\begin{bmatrix}\frac{2p_{m,l}^{0}}{\nu_{l}^{0}}\\ \frac{2q_{m,l}^{0}}{\nu_{l}^{0}}\\ -\frac{(p_{m,l}^{0})^{2}+(q_{m,l}^{0})^{2}}{(\nu_{l}^{0})^{2}}\end{bmatrix}$. (8)
Additionally, we limit squared voltages and currents
@@
Let $\mathcal{S}\subseteq\mathcal{N}$ denote the set of buses equipped with storage units. Each storage unit has a limited capacity $e_{m}^{s}(k)$
@@
with bounded charge/discharge rate
@@
For the sake of simplicity, we neglect the reactive power support of storage systems as well as the charging and discharging efficiency of the storage units.
A simple discrete-time integrator dynamic models the State of Charge (SOC) of the storage units. Assuming a fixed timestep $\Delta k$ and neglecting losses, the SOC dynamic reads
$e_{m}(k+1)=e_{m}(k)-\Delta k\cdot p_{m}^{s}(k),\quad e_{m}(1)=e_{m}^{0}$ (12)
with $m\in\mathcal{S}$. We define the resulting constraint set as
@@
The resulting constraint set is a convex and bounded polyhedron, i.e. a polytope. The set projection of $\mathcal{X}\subseteq\mathcal{Y}\times\mathcal{Z}$ onto $\mathcal{Z}$ is defined as in [undefy]
$\operatorname{proj}_{\mathcal{Z}}(\mathcal{X})\doteq\{z\in\mathcal{Z}\ |\ \exists y\in\mathcal{Y}\text{ with }(z,y)\in\mathcal{X}\}\subseteq\mathcal{Z}.$ (13)
This gives the values for the coupling variables achievable without violating any of the constraints of the original set. Put differently, we can compute the FOR $\mathcal{F}_{i}$ as
@@ III-A Approximation of Feasible Operating Region
with a constraint matrix $A\in\mathbb{R}^{m\times n}$ and a constraint vector $b\in\mathbb{R}^{m}$. This formulation enables the projection of the feasible set onto coupling variables using methods from computational geometry, such as Fourier–Motzkin elimination or vertex enumeration [undefz, pp. 2-3]. However, for multi-stage flexibility aggregation with storage units, the increased problem dimension renders these algorithms computationally intractable.
@@ III-B Reformulation of the Feasible Set
with a center vector $c\in\mathbb{R}^{n}$, a generator matrix $G\in\mathbb{R}^{n\times n_{g}}$ and the scalar internal factors $\alpha_{i}$. Zonotope projection is performed via a simple sparse linear map and therefore can be computed very efficiently. In the following, the shorthand notation $\mathcal{Z}=\langle c,G\rangle_{\mathcal{Z}}$ is used for zonotopes. However, zonotopes are limited to model centrally symmetric convex polytopes. To address this limitation Constrained Zonotopes ($\mathcal{CZ}\mathrm{s}$) extend zonotopes by adding linear constraints to the internal factors $\alpha_{i}$ [undefs]. This results in the definition
@@
with center vector $c\in\mathbb{R}^{n}$, generator matrix $G\in\mathbb{R}^{n\times n_{g}}$, constraint matrix $A\in\mathbb{R}^{m\times n_{g}}$, and constraint vector $b\in\mathbb{R}^{m}$. In the following, the shorthand notation $\mathcal{CZ}=\langle c,G,A,b\rangle_{\mathcal{CZ}}$ is used for constrained zonotopes. Without limitations on the number of generators $n_{g}$ and the number of constraints $m$, $\mathcal{CZ}\mathrm{s}$ can express arbitrary convex polytopes and offer many of the computational advantages of zonotopes, like projection via matrix multiplication [undefs].
@@
Given a feasible set as convex and bounded polytope one can determine upper and lower bounds for each variable by solving two Linear Programms (LPs) for each variable $x_{i,k}$. This allows to determine the vectors of upper and lower variable bounds $ub_{x}$ and $lb_{x}$, which define the bounding zonotope $Z_{b}$ with center $c$ and generator matrix $G$ that over-approximates the given polytope $Z_{b}=\langle c,G\rangle_{Z}\subseteq\mathcal{P}$ as
@@
However, as shown in Subsection V-B selecting the variable bounds sufficiently large is adequate without loss of accuracy, if for each variable $x_{i,k}$ the approximated variable bounds $\tilde{ub_{x}},\tilde{lb_{x}}$ exceed the variable bounds $\tilde{ub_{x}},\tilde{lb_{x}}$ with
@@
Given the feasible set $\mathcal{X}_{i}$ of the multi-stage OPF as polytope $\mathcal{P}$ in H-rep, it is defined as an intersection of a finite number of halfspaces $H_{i}=\{x\in\mathbb{R}^{n}\mid h^{\top}x\leq\zeta\}$. To reconstruct the polytope as $\mathcal{CZ}$, the bounding zonotope is successively intersected with each halfspace $H_{i}$. Based on the method in [undefaa], each intersection updates the $\mathcal{CZ}$ as
@@
This step assumes that the provided polytope is irredundant, so that no feasibility checks are required before intersection. Otherwise, it is beneficial to keep the representation compact by solving an LP to check whether a halfspace intersects the current $\mathcal{CZ}$ and therefore has to be added to the representation [undefaa]. The proposed approach for reformulation of the feasible set as $\mathcal{CZ}$ is summarized in the offline step in Algorithm 1.
@@
1:for all $H_{i}$ modeling static constraints (1), (2), and (5)-(11) do
@@
1:for all $H_{i}$ modeling dynamic constraints like (3), (4), and (12) do
@@
5:Compute (20)
Return $\operatorname{proj}(\mathcal{CZ})$
@@
Finally, once the $\mathcal{CZ}$ is constructed, projection onto the desired subspace is performed via the sparse linear map $M_{p}$, yielding
@@
This projection operation utilizes the computational benefits of the $\mathcal{CZ}$ representation, enabling fast and scalable dimensionality reduction via linear map. Thus, resulting in a computationally efficient online step of Algorithm 1. The additional steps of the online computation for dynamic sets are discussed in Subsection V-B.
Unlike polytope projection that obscures details of sub-grids through variable elimination, the $\mathcal{CZ}$ projection maintains potentially confidential grid information inside $A\textrm{ and }b$ of the equality constraints. In their raw form the constraints might reflect local asset characteristics. However, the offline-online separation of the presented method allows for the seamless integration of privacy-preserving transformations in the offline step, i.e. with a non-singular private transformation matrix $T$ [undefab]
$TA\alpha=Tb,$
which masks the underlying structural data without affecting the communicated FOR.
@@ IV Intersection with Additional Linear Constraints
For more dynamic use cases, it is essential to limit the computation cost required to consider changes in the feasible set $\mathcal{X}_{i}$ due to control inputs or changing external conditions. Given the inherent separation of the set reformulation from the projection in the presented method shown in Algorithm 1, most of the computation can be performed offline. Furthermore, we can utilize that $\mathcal{CZ}\mathrm{s}$ allow to calculate halfspace and hyperplane intersections. This enables an efficient online update strategy for increasing problem sizes, as the underlying topology of the problem remains identical per time-step. We emphasize that, for dynamic use cases, the offline computation incorporates only the strictly static constraints modeling static constraints (1), (2), and (5)-(11) in $\mathcal{X}_{i}$. The offline step itself converts $\mathcal{X}_{i}$ into an equivalent $\mathcal{CZ}_{off}$, as elaborated in the previous section. This offline conversion accounts for the majority of the computational burden. During online operation the offline $\mathcal{CZ}_{off}$ serves as the baseline. Time-varying bounds, such as the integration of revised forecast data in (3), changing power demands or updated setpoint constraints, are integrated by adding the respective constraints via intersection to a copy of $\mathcal{CZ}_{off}$. This results in a new $\mathcal{CZ}_{on}$ modeling the feasible set for each iteration. Because $\mathcal{CZ}_{off}$ deliberately omits previous forecast data, there is no need to computationally remove or modify existing constraints. This avoids a repetition of the offline step entirely. The computational efficiency of adding constraints dynamically via intersections is evaluated in Subsection V-B.